OpenStudy (fibonaccichick666):

ODE clarification, Canonical form: Given \(2y'+siny'=x+y^2~~~and~~~y(0)=0\) Find the canonical form.

OpenStudy (fibonaccichick666):

So I got \[y''=\frac{1+2yy'}{2+cosy'}\] as the canonical form. What am I supposed to do with the y(0)=0? My teacher put "\(2y'(0)+siny'(0)=0+y^2(0)\Rightarrow y'(0)=0\)" on the board

OpenStudy (fibonaccichick666):

@vishweshshrimali5 , any good with this, It's not in my book, so I am jsut not sure on the procedure

OpenStudy (fibonaccichick666):

I just don't see how he solved that for y', like clearly zero works, but is it the only possible solution?

OpenStudy (vishweshshrimali5):

Basically he used y'(0) for finding out the exact value of the integration constant C

OpenStudy (fibonaccichick666):

there is no integration constant

OpenStudy (fibonaccichick666):

no integration going on here

OpenStudy (vishweshshrimali5):

See, whenever you solve a differential equation you basically just integrate it. This integration results in integration constant. A simple differential equation has infinite solutions but one with an initial value has only one solution - because C has been specified

OpenStudy (fibonaccichick666):

but this is not solving, it is canonical form

OpenStudy (fibonaccichick666):

Canonical form is prior to any integration

OpenStudy (vishweshshrimali5):

I am not much familiar with canonical form... sorry :( But, @Kainui and @ganeshie8 can help you out... Again sorry

OpenStudy (fibonaccichick666):

it's cool, I'm just wondering if it was a guess, the real issue is the algebra here: \(2y′(0)+siny′(0)=0+y2(0)⇒y′(0)=0\) He had to solve for y'(0), but how do you solve 2y'+siny'=0?

OpenStudy (dan815):

non-linear odes are not nice

OpenStudy (dan815):

u have to linearize this first

OpenStudy (fibonaccichick666):

No, you don't. I'm not solving it. It's just the canonical form

OpenStudy (fibonaccichick666):

ie get the highest derivative by itself on the left, and that's it. He added this bound like thing and never explained it

OpenStudy (fibonaccichick666):

@Kainui

OpenStudy (fibonaccichick666):

if you want my work for canonical, I can type that up

OpenStudy (kainui):

Yeah that's exactly what I was gonna ask for haha. Show me your work and we'll see what we can figure out from there, I have some ideas.

OpenStudy (fibonaccichick666):

ok so like I said before, my final answer is up top, but her is the work for that: \[2y'+siny'=x+y^2\]\[[2y'+siny']'=[x+y^2]'\]\[2y''+y''cosy'=1+2yy'\]\[y''[2+cosy']=1+2yy'\]\[y''=\frac{1+2yy'}{2+cosy'}\]

OpenStudy (fibonaccichick666):

now he did this problem on the board, he then wrote what I quoted above and said the final answer was what I have + y(0)=0 and y'(0)=0

OpenStudy (kainui):

So it was given that y(0)=0 but then showed that y'(0)=0 too? Or was that also given

OpenStudy (fibonaccichick666):

but the issue I really am having is how (other than guessing) he determined that y'(0)=0

OpenStudy (fibonaccichick666):

nope, showed that y'(0)=0

OpenStudy (fibonaccichick666):

he just plugged and chugged into the original given eq

OpenStudy (fibonaccichick666):

This isn't in the book, so I have no reference

OpenStudy (kainui):

The way I did it is weird, so I'm trying to find a better method one sec.

OpenStudy (fibonaccichick666):

ok

OpenStudy (kainui):

Oh I see! Haha it's kinda fun. Ok check it out, take the original equation and plug in y(0)=0

OpenStudy (fibonaccichick666):

right, I get 2y'+siny'=0

OpenStudy (fibonaccichick666):

but I'm missing why x=0 is given

OpenStudy (kainui):

So now solve for y' in terms of itself we get: \[\Large y'(0)=- \frac{1}{2} \sin( y'(0))\] Here we can ask ourselves, sine of what number times a constant is equal to itself?

OpenStudy (fibonaccichick666):

ohhhhh nice

OpenStudy (fibonaccichick666):

ok that makes sense now

OpenStudy (kainui):

You can draw this as a graph if it is not clear and it will make it immediately apparent, look where y=x and y=-1/2sin(x) intersect.

OpenStudy (fibonaccichick666):

right right, you're right

OpenStudy (fibonaccichick666):

ok, he just didn't explain that reasoning on a new topic....

OpenStudy (fibonaccichick666):

what's the phrase, couldn't see the trees for the forest? I think

OpenStudy (fibonaccichick666):

Thanks

OpenStudy (kainui):

Yeah the original way I did it was ridiculous, I just took this limit by plugging itself in infinitely to itself haha. \[\Large y' = - \frac{1}{2} \sin ( - \frac{1}{2} \sin ( - \frac{1}{2} \sin (...)))\]

OpenStudy (fibonaccichick666):

yea, that may be a bit much lol

OpenStudy (kainui):

But since we have the easy way I just thought for laughs I'd share XD

OpenStudy (fibonaccichick666):

:) thanks laughs are always a good thing.